OpenStudy (anonymous):

Math Analysis: Solve for 0

OpenStudy (anonymous):

let y = sinx so your equation now looks like the quadratic: y^2 + y - 1 = 0 Using the quadratic formula, y = (-(1)[plus/minus]sqrt(1^2-4*(1)*(-1))/(2*1) = (-1[plus/minus]sqrt(5))/2 or y = (-1+sqrt(5))/2 or y = (-1-sqrt(5))/2 Once you have these values, stick it in your calculator: arcsin(y) make sure you're in radian mode.